# Maximum-likelihood estimation of the Matérn covariance structure of isotropic spatial random fields on finite, sampled grids

### v1

@@ 1 I n t r o d u c t i o n
What numbers, which statistical notions capture the “essence” of a spatial patch of geophysical data? If it were a stationary, white, Gaussian process, simply, its population mean and variance would be sufficient. However, computed over differing window sizes, or at varying resolution, sample means and variances of geophysical data sets fluctuate non-erratically, hence “whiteness” immediately proves to be an untenable assumption. Reporting summary statistics over changing baselines [[, e.g.,]for Earth and planetary topography]Sharpton+85,Aharonson+98,Grohmann+2010,Rosenburg+2011 effectively subscribes to the data as realizations from a spatially correlated (locally stationary, Gaussian) process. Estimating even just the variance of said field (that is, at a point), with little bias (that is, accurately) and with a reasonable estimation variance (that is, precisely), requires knowledge of the covariance (between pairs of points). We are led to the estimation of (the parameters of) a spatial covariance function, or alternatively and equivalently, of a spectral density function, from sampled data. The first (spatial estimation) is notoriously noisy since it requires finding data pairs at increasing offsets, and computationally time consuming when it entails inverting a covariance matrix [[, see, e.g.,]]Mardia+84,Kitanidis+85,Vecchia88,Grainger+2021. The second (spectral estimation), while typically faster and less noisy, is famously affected by aliasing, finite-field, and edge effects [[, e.g.,]]Stein95,Appourchaux+98a,Appourchaux+98b,Hamilton2009a,Hamilton2009b. Fourier-domain artifacts lead to estimation bias especially for multidimensional data sets, where the nefarious influence of boundary terms dominates the mean-squared error [Dahlhaus & Künsch(1987)].
@@
Common choices for parameterized covariance functions of Gaussian random fields are exponential or squared-exponential forms [[, e.g.,]]Tarantola+84,Montagner86,Gudmundsson+90,Baig+2003,Baig+2004a, defined solely by a variance and a correlation range. Both are special cases of the [Matérn(1960)] class of covariance functions [Guttorp & Gneiting(2006)], with a specific smoothness or mean-squared differentiability [Adler(1981), Christakos(1992)], itself an inversion parameter of interest in a multitude of geophysical application domains such as seismology [[, e.g.,]]Wu+85b,Wu+90,Carpentier+2007,Becker+2007b, seafloor bathymetry and oceanography [Goff & Jordan(1989a), Goff & Jordan(1989b), Goff & Arbic(2010), Sandwell et al.(2022)Sandwell, Goff, Gevorgian, Harper, Kim, Yu, Tozer, Wessel, & Smith], helioseismology [[, e.g.]]Gizon+2004, hydrology [[, e.g.,]]Rodriguez+74,Kitanidis+85, meteorology, and climate science [[, e.g.,]]Handcock+94,Paciorek+2006,Lindgren+2011,North+2011,Sun+2015.
@@
In this paper we present a univariate spectral-domain debiased “Whittle” maximum-likelihood procedure [Simons & Olhede(2013), Guillaumin et al.(2017)Guillaumin, Sykulski, Olhede, Early, & Lilly, Sykulski et al.(2019)Sykulski, Olhede, Guillaumin, Lilly, & Early] that estimates the variance, smoothness, and range of an isotropic Matérn Gaussian process, from sampled spatial data. We show how to obtain unbiased estimates for these size, scale, and shape parameters when the region studied is neither rectangular nor circular, nor completely sampled. We exactly calculate their estimation covariance, correctly blurred for finite field effects and without neglecting commonly omitted wave vector correlation effects, such that the results from differently sized and sampled patches can be compared robustly. We fully develop and comprehensively illustrate our algorithms for fully sampled rectangular data sets.
@@
Our results have widespread implications for the study of geophysical fields, and should be interpreted in the light of our trying to derive “process” from “parameters”, the end goal being to be able to assign likely formation mechanisms and histories for the patches under consideration. Our results should also be relevant for whomever needs to perform spatial interpolation or out-of-sample extension (e.g. via kriging) on geological data [Journel & Huijbregts(1978), Christakos(1992), Cressie(1993), Stein(1999)]. They also carry consequences for machine learning and feature detection [Rasmussen & Williams(2006), Porcu et al.(2024)Porcu, Bevilacqua, Schaback, & Oates]. We present procedural details but also focus on high-level results that have real-world applications. We illustrate our methodology on four geologically and geophysically relevant data sets, assuming stationarity within patches that have been selected via user interpretation.
@@ 2 P R E L I M I N A R I E S
Readers wishing to come to terms with the geological, geophysical, and geodetic definitions of ‘relief’, ‘topography’, or ‘elevation’ are directed to [Lambeck(1988)], [Hofmann-Wellenhof & Moritz(2006)], and [Wieczorek(2015)]. To make the jump from geology and geophysics to statistics, particularly in this context, we first and foremost recommend (re)reading [Goff & Jordan(1988), Goff & Jordan(1989a)], who also discuss anisotropic processes. The material in this section is both an extension and a specialization of the multivariate results of [Simons & Olhede(2013)], which is to be consulted for further details. Here, we use a more explicit notation, adapt some of the normalizations, and make a number of modifications—but most importantly, we restrict our analysis to univariate two-dimensional Cartesian isotropic Gaussian fields.
@@ 2.1 Continuous framework
Here we draw most heavily on Sections 2.1 and 4.1 of [Simons & Olhede(2013)]. Referring furthermore to [Percival & Walden(1993)], [Stein(1999)], and [Vanmarcke(2010)] for additional considerations and terminology, to [Abramowitz & Stegun(1965)] and [Gradshteyn & Ryzhik(2000)] for properties of special functions, we begin by defining the particular quantities of interest in the spatial and the spectral domains.
@@ 2.1.1 Stationarity
A geophysical field ${\mathcal{H}}(\mathbf{x})$ is considered to be a zero-mean, finite-variance, stationary, two-dimensional random field with finite second moments. Under what is known as the [Cramér(1942)] representation, there exists a spectral orthogonal-increment process, $d{\mathcal{H}}(\mathbf{k})$, according to which the spatial field
@@
The integration is over the space containing all wave vectors $\mathbf{k}$. In the case of strict band-limitation or very fast spectral decay we may restrict the computations to the Nyquist plane $\pi,\pi]\times[-\pi,\pi]$. The expectation of $d{\mathcal{H}}(\mathbf{k})$ over many realized fields,
@@
and its variance, in the absence of co-variance between wave vectors, defines a power-spectral density, ${\mathcal{S}}(\mathbf{k})$, in the form of the expectation
@@ 2.1.2 Isotropy
We follow [Vanmarcke(2010)] in adopting the term ‘fluctuation scale’ for ${\mathcal{S}}(0)/{\sigma^{2}}$, the mass of the spatial correlation function ${\mathcal{C}}(r)/{\sigma^{2}}$.
Isotropy remains a restrictive—but testable—assumption, which we will be relaxing in future work. Any future discussion of anisotropy will entail evaluating it against the null-hypothesis of isotropic behavior, in the possible presence of anisotropic sampling patterns, for which the present paper provides all the necessary statistical machinery.
@@ 2.1.3 Matérnity
We further specify the field as a member of the Matérn class [Stein(1999)], which is very general and widely applicable to processes with monotonically decreasing autocovariance [Guttorp & Gneiting(2006)]. The isotropic $d$-dimensional Matérn spectral density ${\mathcal{S}}^{d}_{\boldsymbol{\theta}}(k)$ assumes the parameterized form [Handcock & Wallis(1994), Wang et al.(2023)Wang, Abdulah, Sun, & Genton]
@@
The ‘variance’, $\sigma^{2}$}, indeed satisfies eq. (9) upon substitution with eq. (12). At short wavelengths, when $k$ is large, the spectrum ${\mathcal{S}}_{\boldsymbol{\theta}}(k)$ decays at a rate that depends on the ‘smoothness’, $\nu$, which expresses the $\lceil\nu-1\rceil$ times (mean-squared) ‘differentiability’ of the process [Handcock & Stein(1993)]. The behavior at the longest wavelengths, for small $k$, is controlled by the combined effect of $\sigma^{2}$} and $\rho$. The fluctuation scale
@@
The isotropic Matérn spatial covariance ${\mathcal{C}}_{\boldsymbol{\theta}}(r)$, in terms of the lag distance $r$, is unlike its spectral counterpart (11) in requiring no dimensional specification, that is, independent of $d$,
@@
The flexible generality of the Matérn class is appreciated by evaluating the correlation functions for special values of $\nu$ [Guttorp & Gneiting(2006)]. Notably, when $\nu=1/2$, the correlation function decays exponentially, and when $\nu\rightarrow\infty$, as a Gaussian—a squared exponential. Other examples include the Von Kármán ($\nu=1/3$), Whittle ($\nu=1$), and second-order ($\nu=3/2$) and third-order ($\nu=5/2$) autoregressive correlation models. Despite all of its generality and wide applicability, it is important to not identify our conceptual point of departure with (multi-)fractal, scale-invariant, self-affine, or self-similar behavior [[, see, e.g.,]]Mareschal89,Herzfeld+95,Gneiting+2012,Landais+2019, about which we make no claims.
Fig. 1 provides intuitive insight into the role that the three parameters $\sigma^{2}$}, $\nu$, and $\rho$ play in the spatial behavior of Matérn random fields, synthesized by the procedure outlined in the next section.
Figure 1: Random fields generated from stationary isotropic Matérn models with variances $\sigma^{2}$}, differentiabilities $\nu$, and correlation lengths $\rho$, as indicated. (Top:) Normalized spectral densities, ${\mathcal{S}}_{\boldsymbol{\theta}}(k)/{\mathcal{S}}_{\boldsymbol{\theta}}(0)$, from eq. (12). The vertical black lines identify the wavenumbers $k_{\alpha}$ at which the power reaches $100\times\alpha$ per cent of the variance, from eq. (18), in wavelengths $\lambda_{100\alpha}=2\pi/k_{\alpha}$, as labeled. (Middle:) Correlations, the normalized spatial covariances, ${\mathcal{C}}_{\boldsymbol{\theta}}(r)/{\sigma^{2}}$, from eq. (15). The vertical blue lines are drawn at the values $\pi\rho$, the distances at which the correlations die down to approximately one third of the variance (Bottom:) Field realizations. The blue circles have radii $\pi\rho$, drawn for visual guidance. In the titles, $\mathsf{m}$ and $\mathsf{s}$ identify the sample means and standard deviations.
@@ 2.2 Lattice framework
Here we rely mostly on sections 2.1, 4.2, and A6 of [Simons & Olhede(2013)]. The properties of the finite and sampled, i.e., windowed discrete processes, as will be experienced in computational data analysis, differ markedly from the behavior of the idealized, infinite, continuous models discussed in the previous section, and those two viewpoints need to be explicitly reconciled.
@@ 2.2.1 Discretization
Sampled in spectral space, the finite set wave vectors is now, with $m$ and $n$ as in eq. (19),
@@ 2.2.2 Blurring
[Simons & Olhede(2013)] showed the covariance of a finite set of gridded Fourier coefficients can at best offer a blurred and correlated version of the true spectral variance [[, see also]]Fournier+2014. To us, their eq. (9) reads, using ${\mathcal{S}}_{\boldsymbol{\theta}}(k)={\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$ from our eq. (12),
$\langle H(\mathbf{k})H^{*}\hskip-1.00006pt(\mathbf{k^{\prime}})\rangle=\int\!\!\!\!\int D_{K}(\mathbf{k}-\mathbf{k}^{\prime\prime})D_{K}^{*}(\mathbf{k}^{\prime}-\mathbf{k}^{\prime\prime})\hskip 1.00006pt{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k}^{\prime\prime})\,d\mathbf{k}^{\prime\prime}\quad\mbox{with}\quad D_{K}(\mathbf{k})=\frac{1}{2\pi}\left(\frac{\Delta x\Delta y}{MN}\right)^{\frac{1}{2}}\sum_{\mathbf{x}}e^{-i\mathbf{k}\cdot\mathbf{x}},$ (23)
@@
$\langle H(\mathbf{k})H^{*}\hskip-1.00006pt(\mathbf{k^{\prime}})\rangle\approx\delta(\mathbf{k},\mathbf{k}^{\prime})\int\!\!\!\!\int\left|D_{K}(\mathbf{k}-\mathbf{k}^{\prime})\right|^{2}{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k^{\prime}})\,d\mathbf{k}^{\prime}=\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})\delta(\mathbf{k},\mathbf{k}^{\prime}).$ (24)
Furthermore, they carried out the blurring computationally, and approximately, via grid refinement, convolution, and subsampling.
[Guillaumin et al.(2022)Guillaumin, Sykulski, Olhede, & Simons], in contrast, in their Lemmata 1 and 2, showed how to exactly incorporate the spectral blurring effect of applying arbitrary data windows [[, see also]]Fuentes2007, including irregular boundaries and incomplete sampling, at a much reduced computational cost. We rewrite the discrete Fourier transform in eq. (20) to incorporate an arbitrary unit-normalized data window, $w(\mathbf{x})$, as follows:
@@
Eq. (30) is the exact version, valid for arbitrary data windows, of what [Simons & Olhede(2013)] implemented approximately and only for rectangular windows by discrete convolution of the theoretical spectral density ${\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$ with the Fejér kernel $\left|D_{K}(\mathbf{k})\right|^{2}$, as in eq. (24). In that special case of a unitary window function, that is, for complete observations on a rectangular grid, eq. (29) evaluates to the triangular
@@
Used as a basis for simulation, generating Fourier coefficients $H(\mathbf{k})$ from the square root of the blurred spectral density $\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$ of eq. (30) ignores wavenumber correlation effects. To make Fig. 1, we generated samples on a spatial grid four times as large as needed, retaining only the central portion for analysis, to avoid wrap-around correlations or periodic realizations. Constructing spatial patches via eq. (22), on the space grid (19), the variance of the results, $\langle{\mathcal{H}}(\mathbf{x}){\mathcal{H}}\hskip-1.00006pt(\mathbf{x})\rangle$, may be understood as a discrete approximation of the integral in eq. (4), with the spacings defined in eq. (21). Hermitian symmetry guarantees that the simulated fields are real, and their covariance
@@
Eq. (30) shows that the expected periodogram of the data, $\langle|H(\mathbf{k})|^{2}\rangle$, can be obtained via Fourier transformation of the autocovariance sequence of the sampling window. Fig. 2 shows this equivalence. From a sequence of realizations, we show, in the top row from left to right, one spatial-domain field, ${\mathcal{H}}(\mathbf{x})$, its periodogram $|H(\mathbf{k})|^{2}$ on the corresponding normalized Fourier grid, and the expected periodogram, the blurred spectrum $\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$, for the parameter set $\boldsymbol{\theta}$ shown at the top, ${\sigma^{2}}=1$ (in arbitrary field units), $\nu=2.5$ and $\rho=1$ (in units of the spatial grid spacing). In the bottom row, we show a unit normalized square taper, the ratio of the average periodogram to its expectation, $\mathrm{mean}\{|H(\mathbf{k})|^{2}\}/\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$, with its sample mean $m$ and standard deviation $s$, and the average of the periodograms across 100 realizations, as the sample variance, $\mbox{var}\{H(\mathbf{k})\}$, which approximates the blurred spectrum $\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$ shown directly above.
Figure 2: The Fejér blurred spectral density $\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$ approximates the expectation of the periodogram, $|H(\mathbf{k})|^{2}$, of gridded and (unit) tapered data generated from the population density ${\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$. (Top row) A single realization, ${\mathcal{H}}(\mathbf{x})$, its modified periodogram $|H(\mathbf{k})|^{2}$, and the blurred spectrum $\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$. (Bottom row) The unit taper, the ratio of the average periodogram to the blurred spectral density, and the average periodogram, over 100 realizations.

### v2

@@ 1 I n t r o d u c t i o n
What numbers, which statistical notions capture the “essence” of a spatial patch of geophysical data? If it were a stationary, white, Gaussian process, simply, its population mean and variance would be sufficient. However, computed over differing window sizes, or at varying resolution, sample means and variances of geophysical data sets fluctuate non-erratically, hence “whiteness” immediately proves to be an untenable assumption. Reporting summary statistics over changing baselines [[, e.g.,]for Earth and planetary topography]Sharpton+85,Aharonson+98,Grohmann+2010,Rosenburg+2011 effectively subscribes to the data as realizations from a spatially correlated (locally stationary, Gaussian) process. Estimating even just the variance of said field (that is, at a point), with little bias (that is, accurately) and with a reasonable estimation variance (that is, precisely), requires knowledge of the co-variance (between pairs of points), and an acknowledgment of the inherently finite nature of the available spatial patch over which a single (geological) process can be thought of as representative. We are led to the estimation of (the parameters of) a spatial covariance function, or alternatively and equivalently, of a spectral density function, from (hopefully locally) stationary sampled data that are (irregularly) bounded [[, see, e.g.,]]Valentine+2020. The first (spatial estimation) is notoriously noisy since it requires finding data pairs at increasing offsets, and computationally time-consuming when it entails inverting a covariance matrix [[, see, e.g.,]]Mardia+84,Kitanidis+85,Vecchia88,Grainger+2021. The second (spectral estimation), while typically faster and less noisy, is famously affected by aliasing, finite-field, and edge effects [[, e.g.,]]Stein95,Appourchaux+98a,Appourchaux+98b,Hamilton2009a,Hamilton2009b. Fourier-domain artifacts lead to estimation bias especially for multidimensional data sets, where the nefarious influence of boundary terms dominates the mean-squared error [11].
Estimating model parameters of random fields typically relies on spatial [40] or spectral [68] likelihood maximization. If the former is often slow to compute, and the latter suffers from potential boundary effects, there are notable modifications to both. For the spatial methods, Vecchia approximations to Gaussian log-likelihoods are fast and accurate [[, see, e.g.,]]Katzfuss+2021,Porcu+2024. Circulant-embedding based methodologies [29, 30] effectively recompute the Fourier basis rather than assuming it known from homogeneity in space. Vecchia-type approaches do require a local neighborhood for prediction, chosen appropriately. Some of these methods are discussed relative to spectral approaches both in theory and via simulations by [28], who furthermore introduce adjustments to correct for partial-grid and boundary effects with the spectral methods, developing a large-sample theory for the appropriate asymptotic framework, upon which we build in this paper.
@@
Common choices for parameterized covariance functions of Gaussian random fields are exponential or squared-exponential forms [[, e.g.,]]Tarantola+84,Montagner86,Gudmundsson+90,Baig+2003,Baig+2004a, defined solely by a variance and a correlation range. Both are special cases of the [44] class of covariance functions [31], with a specific smoothness or mean-squared differentiability [2, 6], itself an inversion parameter of interest in a multitude of geophysical application domains such as seismology [[, e.g.,]]Wu+85b,Wu+90,Becker+2007b,Carpentier+2007, seafloor bathymetry and oceanography [22, 21, 19, 55, 57], helioseismology [[, e.g.,]]Gizon+2004, hydrology [[, e.g.,]]Rodriguez+74,Kitanidis+85, meteorology, and climate science [[, e.g.,]]Handcock+94,Paciorek+2006,Lindgren+2011,North+2011,Sun+2015.
@@
In this paper we present a univariate spectral-domain debiased “Whittle” maximum-likelihood procedure [58, 26, 62] that estimates the variance, smoothness, and range of an isotropic Matérn Gaussian process, from sampled spatial data. We show how to obtain unbiased estimates for these size, scale, and shape parameters when the region studied is neither rectangular nor circular, nor completely sampled. We exactly calculate their estimation covariance, correctly blurred for finite-field effects and without neglecting commonly omitted wave vector correlation effects, such that the results from differently sized and sampled patches can be compared robustly. We develop and comprehensively illustrate our algorithms for completely sampled rectangular data sets.
While process variance estimation necessitates recognizing the effects of spatial covariance, which is aided by the ability to extend the size of the observation domain, intuitively “beyond the reach of the range”, conversely, to characterize the smoothness of a process requires neighboring observations that are highly correlated. In other words, both increasing (growing) domain and fixed domain (infill) asymptotics [[, and mixed-domain, see]]Chu2023 are relevant considerations [[, see, e.g.,]]Zhang2004,Zhang+2005b that we are, however, unable to fully address in this paper, although the computer code that we freely distribute (see Sec. 7) is designed to give the reader all the tools in hand to be able to do so. Neither form of asymptotics is a panacea, and, furthermore, an unfortunate aspect of the Matérn class is that in two dimensions, neither the variance nor the range are consistently estimable under fixed-domain asymptotics, although there is a non-linear function of them and the smoothness that is [[, see]]Stein95. [28], who provided the theoretical basis for our work, introduced the notion of significant correlation contribution as a means of capturing whether a covariance function under a given spatial sampling mechanism allows us to distinguish parameter values. It cannot eliminate intrinsic problems with the flexibility of the Matérn family, and in some scenarios it may indeed be beneficial to fix the smoothness. While the full theoretical framework allows us to determine when we are able to estimate a specific Matérn family with a given sampling mechanism, it is our goal, with this paper, to provide sufficient, but not exhaustive, guidance, accompanied by computer code, to make Matérn covariance estimation practical and sensible for geoscientists.
@@
Our results have widespread implications for the study of geophysical fields, and should be interpreted in the light of our trying to derive “process” from “parameters”, the end goal being to be able to assign likely formation mechanisms and histories for the patches under consideration. Our results should also be relevant for whomever needs to perform spatial interpolation or out-of-sample extension (e.g., via kriging) on geological data [39, 6, 9, 61]. They also carry consequences for machine learning and feature detection [53, 50]. We present procedural details but also focus on high-level results that have real-world applications. We illustrate our methodology on four geologically and geophysically relevant data sets, assuming stationarity within patches that have been selected via user interpretation.
@@ 2 P R E L I M I N A R I E S
Readers wishing to come to terms with the geological, geophysical, and geodetic definitions of ‘relief’, ‘topography’, or ‘elevation’ are directed to [42], [36], and [70]. To make the jump from geology and geophysics to statistics, particularly in this context, we first and foremost recommend (re)reading [20, 22], who also discuss anisotropic processes. The material in this section is both an extension and a specialization of the multivariate results of [58], which is to be consulted for further details. Here, we use a more explicit notation, adapt some of the normalizations, and make a number of modifications—but most importantly, we restrict our analysis to univariate two-dimensional Cartesian isotropic Gaussian fields.
@@ 2.1 Continuous framework
Here we draw most heavily on Sections 2.1 and 4.1 of [58]. Referring furthermore to [49], [61], and [64] for additional considerations and terminology, to [1] and [23] for properties of special functions, we begin by defining the particular quantities of interest in the spatial and the spectral domains.
@@ 2.1.1 Stationarity
A geophysical field ${\mathcal{H}}(\mathbf{x})$ is considered to be a zero-mean, finite-variance, stationary, two-dimensional random field with finite second moments. Under what is known as the [8] representation, there exists a spectral orthogonal-increment process, $d{\mathcal{H}}(\mathbf{k})$, according to which the spatial field
@@
The integration is over the space containing all wave vectors $\mathbf{k}$. In the case of strict bandlimitation or very fast spectral decay we may restrict the computations to the Nyquist plane $\pi,\pi]\times[-\pi,\pi]$. The expectation of $d{\mathcal{H}}(\mathbf{k})$ over many realized fields,
@@
and its variance, in the absence of covariance between wave vectors, defines a power-spectral density, ${\mathcal{S}}(\mathbf{k})$, in the form of the expectation
@@ 2.1.2 Isotropy
We follow [64] in adopting the term ‘fluctuation scale’ for ${\mathcal{S}}(0)/{\sigma^{2}}$, the mass of the spatial correlation function ${\mathcal{C}}(r)/{\sigma^{2}}$.
Isotropy remains a restrictive—but testable—assumption, which we will be relaxing in forthcoming work. Any future discussion of anisotropy will entail evaluating it against the null-hypothesis of isotropic behavior, in the possible presence of anisotropic sampling patterns, for which the present paper provides all the necessary statistical machinery.
@@ 2.1.3 Matérnity
We further specify the field as a member of the Matérn class [61], which is very general and widely applicable to processes with monotonically decreasing autocovariance [31]. The isotropic $d$-dimensional Matérn spectral density ${\mathcal{S}}^{d}_{\boldsymbol{\theta}}(k)$ assumes the parameterized form [34, 66]
@@
The ‘variance’, $\sigma^{2}$}, indeed satisfies eq. (9) upon substitution with eq. (12). At short wavelengths, when $k$ is large, the spectrum ${\mathcal{S}}_{\boldsymbol{\theta}}(k)$ decays at a rate that depends on the ‘smoothness’, $\nu$, which expresses the $\lceil\nu-1\rceil$ times (mean-squared) ‘differentiability’ of the process [33]. The behavior at the longest wavelengths, for small $k$, is controlled by the combined effect of $\sigma^{2}$} and $\rho$. The fluctuation scale
@@
The isotropic Matérn spatial covariance ${\mathcal{C}}_{\boldsymbol{\theta}}(r)$, in terms of the lag distance $r$, is unlike its spectral counterpart (11) in requiring no dimensional specification, that is, independently of $d$,
@@
The flexible generality of the Matérn class is appreciated by evaluating the correlation functions for special values of $\nu$ [31]. Notably, when $\nu=1/2$, the correlation function decays exponentially, and when $\nu\rightarrow\infty$, as a Gaussian—a squared exponential. Other examples include the Von Kármán ($\nu=1/3$), Whittle ($\nu=1$), and second-order ($\nu=3/2$) and third-order ($\nu=5/2$) autoregressive correlation models. Despite all of its generality and wide applicability, the exponentially decaying, isotropic, three-parameter Matérn model cannot capture all random fields under study. It is not identifiable with (multi-) fractal, scale-invariant, self-affine, or self-similar behavior [[, see, e.g.,]]Mareschal89,Herzfeld+95,Gneiting+2012,Landais+2019, as it is self-affine only at high wavenumber. For temporal processes, [43] discuss when the Matérn process is more appropriate to use than fractional Brownian motion. Developing a general set of methods outside these constraints, e.g., when the spectrum is no longer bounded [54], remains outstanding as future work.
Fig. 1 provides intuitive insight into the role that the three parameters $\sigma^{2}$}, $\nu$, and $\rho$ play in the spatial behavior of Matérn random fields, synthesized by the circulant embedding procedure outlined in the next section.
Figure 1: Random fields generated from stationary isotropic Matérn models with variances $\sigma^{2}$}, differentiabilities $\nu$, and correlation lengths $\rho$, as indicated. (Top:) Normalized spectral densities, ${\mathcal{S}}_{\boldsymbol{\theta}}(k)/{\mathcal{S}}_{\boldsymbol{\theta}}(0)$, from eq. (12). The vertical black lines identify the wavenumbers $k_{\alpha}$ at which the power reaches $100\times\alpha$ per cent of the variance, from eq. (18), in wavelengths $\lambda_{100\alpha}=2\pi/k_{\alpha}$, as labeled. (Middle:) Correlations, the normalized spatial covariances, ${\mathcal{C}}_{\boldsymbol{\theta}}(r)/{\sigma^{2}}$, from eq. (15). The vertical blue lines are drawn at the values $\pi\rho$, the distances at which the correlations die down to approximately one third of the variance. (Bottom:) Field realizations. The blue circles have radii $\pi\rho$, drawn for visual guidance. In the titles, $\mathsf{m}$ and $\mathsf{s}$ identify the sample means and standard deviations.
@@ 2.2 Lattice framework
Here we rely mostly on sections 2.1, 4.2, and A6 of [58]. The properties of the finite and sampled, i.e., windowed discrete processes, as will be experienced in computational data analysis, differ markedly from the behavior of the idealized, infinite, continuous models discussed in the previous section, and those two viewpoints need to be explicitly reconciled.
@@ 2.2.1 Discretization
Sampled in spectral space, the finite set of wave vectors is now, with $m$ and $n$ as in eq. (19),
@@ 2.2.2 Blurring
[58] showed the covariance of a finite set of gridded Fourier coefficients can at best offer a blurred and correlated version of the true spectral variance [[, see also]]Fournier+2014. To us, their eq. (9) reads, using ${\mathcal{S}}_{\boldsymbol{\theta}}(k)={\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$ from our eq. (12),
$\langle H(\mathbf{k})H^{*}\hskip-1.00006pt(\mathbf{k^{\prime}})\rangle=\iint\limits_{\mathbf{k}}D_{K}(\mathbf{k}-\mathbf{k}^{\prime\prime})D_{K}^{*}(\mathbf{k}^{\prime}-\mathbf{k}^{\prime\prime})\hskip 1.00006pt{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k}^{\prime\prime})\,d\mathbf{k}^{\prime\prime}\quad\mbox{with}\quad D_{K}(\mathbf{k})=\frac{1}{2\pi}\left(\frac{\Delta x\Delta y}{MN}\right)^{\frac{1}{2}}\sum_{\mathbf{x}}e^{-i\mathbf{k}\cdot\mathbf{x}},$ (23)
@@
$\langle H(\mathbf{k})H^{*}\hskip-1.00006pt(\mathbf{k^{\prime}})\rangle\approx\delta(\mathbf{k},\mathbf{k}^{\prime})\iint\limits_{\mathbf{k}}\left|D_{K}(\mathbf{k}-\mathbf{k}^{\prime})\right|^{2}{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k^{\prime}})\,d\mathbf{k}^{\prime}=\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})\delta(\mathbf{k},\mathbf{k}^{\prime}).$ (24)
The blurred spectral density $\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$ is the convolution of the true spectral density ${\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$ with the Fejér kernel $\left|D_{K}(\mathbf{k}-\mathbf{k}^{\prime})\right|^{2}$, which embodies the effects both of sampling and finite domain size, and reduces to the Dirac delta for continuous processes on infinite domains. [58] carried out the blurring approximately, via grid refinement, convolution, and subsampling.
[28], in contrast, in their Lemmata 1 and 2, showed how to exactly incorporate the spectral blurring effect of applying arbitrary data windows [[, see also]]Fuentes2007, including irregular boundaries and incomplete sampling, at a much reduced computational cost. We rewrite the discrete Fourier transform in eq. (20), which silently selected samples on the grid $\mathbf{x}$ but did not modify or weight them, to incorporate a non-constant, unit-normalized, data window, $w(\mathbf{x})$, as follows:
@@
Eq. (30) is the exact version, valid for arbitrary data windows, of what [58] implemented approximately and only for rectangular windows by discrete convolution of the theoretical spectral density ${\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$ with the Fejér kernel $\left|D_{K}(\mathbf{k})\right|^{2}$, as in eq. (24). In that special case of a unitary window function, that is, for complete observations on a rectangular grid, eq. (29) evaluates to the triangular
@@
Generating Fourier coefficients $H(\mathbf{k})$ from the square root of the blurred spectral density $\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$ of eq. (30), as [58] did as a basis for simulation, ignores wavenumber correlation effects. In that case samples need to be generated on a spatial grid larger than needed, retaining only a central portion for analysis, to avoid wrap-around correlations or periodic realizations. Constructing spatial patches via eq. (22) on the space grid (19), the covariance of the results, $\langle{\mathcal{H}}(\mathbf{x}){\mathcal{H}}\hskip-1.00006pt(\mathbf{x}^{\prime})\rangle$, may be understood as a discrete approximation of the integral in eq. (4), with the spacings defined in eq. (21). Hermitian symmetry guarantees that the simulated fields are real, and their covariance
@@
Eq. (30) shows that the expected periodogram of the data, $\mbox{var}\{H(\mathbf{k})\}=\langle|H(\mathbf{k})|^{2}\rangle$, can be obtained via Fourier transformation of the autocovariance sequence of the sampling window. Fig. 2 shows this equivalence. From a sequence of simulations, we show, in Fig. 2(a–c), one spatial-domain field ${\mathcal{H}}(\mathbf{x})$, its periodogram $|H(\mathbf{k})|^{2}$ on the corresponding normalized Fourier grid, and the expected periodogram, the blurred spectrum $\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$, for the parameter set $\boldsymbol{\theta}$ shown at the top, ${\sigma^{2}}=1$ (in arbitrary field units), $\nu=2.5$ and $\rho=1$ (in units of the spatial grid spacing). In Fig. 2(d–f), we show a unit-normalized square window cosine-tapered around the edges, the ratio of the average periodogram to its expectation, $\mathrm{mean}\{|H(\mathbf{k})|^{2}\}/\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$, with its sample mean $\mathsf{m}$ and standard deviation $\mathsf{s}$, and the average of the periodograms across 100 realizations, the sample variance $\mbox{var}\{H(\mathbf{k})\}$, which approximates the blurred spectrum $\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$ shown directly above.
Figure 2: The Fejér blurred spectral density $\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$ approximates the expectation of the periodogram $|H(\mathbf{k})|^{2}$, of gridded and cosine-tapered data generated from the population density ${\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$. (a) A single realization ${\mathcal{H}}(\mathbf{x})$, (b) its modified periodogram $|H(\mathbf{k})|^{2}$, and (c) the blurred spectrum $\bar{\mathcal{S}}_{\boldsymbol{\theta}}(\mathbf{k})$. (d) The cosine-tapered unit window, (e) the ratio of the average periodogram to the blurred spectral density, and (f) the average periodogram, over 100 realizations.